1:The average amount of miles a car battery lasts for in a 2019 sedan is 37,452 miles with a standard deviation of

4,890 miles. If this distribution is normally distributed, answer the following questions:

A: What is the probability you select a battery it last longer than 40,000 miles?

B: What is the probability you select a battery, and it lasts between 32,000 miles and 44,000 miles?

C: Where does the top 18% of longest lasting batteries begin (at what mileage)?


1
Expert's answer
2022-03-11T01:42:46-0500

Mean "\mu = 37,452"

Standard deviation "\sigma =4,890"


(a) Battery last longer than "40,000"

"Z=\dfrac{X-\mu}{\sigma}"


"Z=\dfrac{40,000-37,452}{4,890}=0.5211"


From normal distribution tables

P(Z of 0.5211) "=0.69847"


The battery lasting longer

"1-0.69846=0.30154"


(b) Battery lasting between "32,000" and "44,000"

"Z=\dfrac{X-\mu}{\sigma}"


Z1 "=\dfrac{32,000-37,452}{4,890}=-1.115"


Z2 "=\dfrac{44,000-37,452}{4,890}=1.339"


From normal distribution tables

P1 "=0.1335"

P2 "= 1-0.90988 =0.09012"


"P(32,000\to44,000)=1-(0.1335+0.09012)"

"=0.77638"


(c) Where does top "18\%" begin

Top "18\%" begins at "100-18=82\%"


"P=0.82"

From normal distribution tables

Z(for P=0.82) "=0.92"


"Z=\dfrac{X-\mu}{\sigma}"


"X=σZ+\mu"

"X=0.92\times4,890+37,452"

"X=41,590" miles